# Deep Discrete Encoders: Identifiable Deep Generative Models for Rich Data with Discrete Latent Layers

### v1

@@ 5 Simulation Studies
We conduct extensive simulation studies in various settings to assess the performance of the proposed computation pipeline (Algorithms 2 and 3) and validate our identifiability conditions (in Section 3.1). We generate data from two-latent-layer DDEs, exploring a total of 90 true settings by varying the following:
@@
    two parameter values: see (13) below,
@@
Given a value of $K^{(2)}$, set $K^{(1)}=3K^{(2)}$ and $J=9K^{(2)}$, which allows a large latent dimension with as many as $K^{(1)}=30$ binary latent variables in the shallowest latent layer.
In terms of the true parameter values, we consider two sets of values based on the strict and generic identifiability conditions in Theorems 1 and 2, respectively. In particular, we define $\mathcal{B}_{s}=\{\mathbf{B}_{s}^{(d)}\}_{d=1,2}$ that satisfy Theorem 1 as follows:
$\mathbf{B}_{s}^{(d)}=\begin{pmatrix}-2\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -4\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{1}^{(d)}\end{pmatrix},\quad\text{where}\quad\beta_{1;j,k}^{(d)}:=\begin{cases}4&\text{if }k=j,\\ 4/3&\text{ else if }|k-j|=K^{(d)}/2,\\ 0&\text{otherwise}.\end{cases}$ (13)
We also consider $\mathcal{B}_{g}=\{\mathbf{B}_{g}^{(d)}\}_{d=1,2}$ that satisfy Theorem 2, whose exact values are displayed in Supplement S.4.1. We set the top-layer proportion parameters as $p_{k}=0.5$ for all $k\in[K^{(2)}]$, and set the dispersion parameters for the Normal distribution as $\gamma_{j}=\sigma_{j}^{2}=1$ for all $j\in[J]$.
@@
Figure 3: Estimation error for $\mathcal{G}$ and $\boldsymbol{\Theta}$ under the two-latent-layer DDE with parameters $\mathcal{B}_{s}$ and various observed-layer parametric families.
@@
For each scenario, we run $100$ independent simulations, and display the estimation results in Figure 3. The results under the generic identifiable parameters $\mathcal{B}_{g}$ and computation times are included in Supplementary Material S.4.1. We measure the estimation accuracy of the graphical matrices $\mathcal{G}$ by computing the average entrywise accuracy. For the continuous parameters $\boldsymbol{\Theta}$, we report the root mean squared error (RMSE). Under all response types and parameter values, the estimation errors for both $\mathcal{G}$ and $\boldsymbol{\Theta}$ decrease as the sample size $N$ increases. This empirically justifies the identifiability and consistency results. Additionally, by comparing the estimation accuracy across different parametric families, we observe that the Bernoulli is the most challenging to estimate, and the Normal is the easiest. Regarding the true parameters, the sparse model with coefficients $\mathcal{B}_{s}$ is easier to estimate. This is because sparsity helps the initialization and reduces the effective parameter dimension. Finally, under coefficients $\mathcal{B}_{s}$, estimation accuracy improves as the latent and observed dimensions increase, which is a consequence of the decreasing percentage of nonzero entries in $\mathcal{B}$.
@@
In Supplementary Material S.4.3, we present simulation results where the latent dimension $\mathcal{K}$ is unknown. We illustrate that the spectral-gap estimator has near-perfect selection accuracy for a large $N$ (e.g., larger than $4000$), and is superior in terms of both accuracy and computation time compared to alternative estimation strategies.
@@ 6.1 Binary Data: Bernoulli-DDE for MNIST Handwritten Digits
Table 1: From left to right: Estimated basis images ${\boldsymbol{\beta}}_{0,0}^{(1)}$, ${\boldsymbol{\beta}}_{0,1}^{(1)}$, ${\boldsymbol{\beta}}_{0,2}^{(1)}$, ${\boldsymbol{\beta}}_{0,3}^{(1)}$, ${\boldsymbol{\beta}}_{0,4}^{(1)}$, ${\boldsymbol{\beta}}_{0,5}^{(1)}$ from the MNIST data, which define sparse subregions in the $28\times 28$ image. The second and third row shows the most significant portion of the basis images by thresholding at the value ±$\pm 1.5$.
We fit the two-latent-layer DDE using Algorithm 1 with the latent dimensions set to $K^{(1)}=5$ and $K^{(2)}=2$ (see the Supplementary Material S.5.2 for the rationale for this choice). Table 1 displays the first-layer coefficients $\mathbf{B}^{(1)}$ by rearranging the each column into the original $28\times 28$ grid. Note that the value of $\mathbf{B}^{(1)}$ is in the logit-scale, so the negative coefficients make the corresponding pixel more likely to be zero. From the reshaped columns ${\boldsymbol{\beta}}_{0,0}^{(1)}$, ${\boldsymbol{\beta}}_{0,1}^{(1)}$, …, ${\boldsymbol{\beta}}_{0,5}^{(1)}$ of $\mathbf{B}^{(1)}$, we can interpret the meaning of each latent variable: $A_{1}^{(1)}=1$ indicates a zero-like shape, $A_{5}^{(1)}=1$ indicates symmetric curves on the upper-left and bottom-right corners, and the other latent variables represent different rotations. Additionally, using the deeper graphical matrix $\mathbf{G}^{(2)}$ displayed in the Supplementary Material S.5.4, we can also interpret the top layer latent variables as broader information about the images. For example, $A_{1}^{(2)}$ indicates large pixel density and $A_{2}^{(2)}$ indicates symmetry with respect to the x-axis.
@@
$(\widehat{\mathbf{A}}^{(1)}_{i},\widehat{\mathbf{A}}^{(2)}_{i})=\operatorname*{argmax}_{\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}},~{}\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K_{2}}}\varphi_{i,\boldsymbol{\alpha}^{(1)},\boldsymbol{\alpha}^{(2)}};\quad\boldsymbol{\alpha}^{(\ell)}\in\{0,1\}^{K_{\ell}},~{}\ell=1,2.$ (14)
Then, a decision tree that classifies the binary latent representations to the categorical label is built by using the misclassification error as the splitting criteria; see the left panel of Figure 4. The center and right panels display the estimated latent traits and estimated labels. Our classifier leads to a high train/test accuracy of $92.0\%/92.6\%$, even though our model is not fine-tuned for image classification. Although the state-of-art machine learning methods can achieve an accuracy as high as $97\%$ (Monnier et al.,, 2020), we point out that the main goal here is not classification, but on interpretability and parsimony; indeed, the DDE uses more limited information and is a less complex but more interpretable model that provides the generative process for the image. In Table 3, we compare the classification accuracy and pixel-wise reconstruction accuracy of the two-latent-layer DDE with alternative interpretable models, such as the latent class model (LCM) and the one-latent-layer DDE. The results show that the two-latent-layer DDE performs the best. We also display example generated images alongside their latent configurations in Table 2, which illustrates various handwriting styles for each digit. Supplementary Material S.5.4 additionally displays sample reconstructed images based on these three methods, and illustrates that the LCM and the one-latent-layer DDE generate noisier and less flexible images than the two-latent-layer DDE.
$[1,0,1,0,1]$ $[0,0,0,0,0]$ $[0,1,0,1,1]$ $[0,0,1,1,1]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $[1,0,1,1,1]$ $[0,0,0,1,0]$ $[0,1,1,0,1]$ $[0,0,1,1,0]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Table 2: Example generated images from the two-latent-layer DDE and their latent representations. Accuracy LCM 1-DDE 2-DDE Train classif. (%) 89.5 84.8 92.0 Test classif. (%) 90.9 86.2 92.6 Train recon. (%) 48.2 79.5 79.6 Test recon. (%) 48.5 79.1 79.9 Table 3: Classification accuracy and pixel-wise reconstruction accuracy for the MNIST dataset.
@@ 6.2 Count Data: Poisson-DDE Hierarchical Topic Modeling
We analyze the text corpus from the 20 newsgroups dataset (Lang,, 1995), which was previously analyzed by other topic models with binary latent variables (Srivastava et al.,, 2013; Gan et al.,, 2015). The training set of the corpus consists of approximately 11,000 documents and 2,000 words, and is based on the email messages across 20 newsgroups. After preprocessing by focusing on 12 newsgroups and removing stop words and infrequent words, the dataset consists of $N=5,883$ documents and $J=653$ words. While not used for fitting the DDE, the label (newsgroup) of each document is also provided in the dataset, alongside three larger categories: recreation, computer, and science.
Y1.0e0]ptrecreationtechnologycarssportsmotorcyclesoftwarelocation/nameshardwaregraphicsspacegeb bike banks riding pitt dod bmw gordonhockeynhlplayoffsteamsplayersleagueseasonbaseballenginecheappartsturnedmodelsdealermarketcarfilesdirectoryfilemscopydeletedsizecharlespeter ny chicagodevicesperiodyorkthomasusamotherboard dx CPU mhz controller board ram cardsmouse motif lines color display screen window imageorbit earth research science nasa development project institute
@@
We fit the two-latent-layer DDE with a Poisson-distributed data layer (Poisson-DDE) $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{Poi}\big{(}\exp\big{[}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)}\big{]}\big{)}$, and display the estimated latent structure in Figure 5. We set $K^{(1)}=8$ based on the spectral ratio estimator and $K^{(2)}=2$ based on the identifiability conditions and better interpretability. Additional details behind this choice are given in Supplementary Material S.5.2. Our choice of $K^{(2)}=2$ is also consistent with the correlation structure discovered in Gan et al., (2015). To better interpret individual latent variables, we define representative words for each topic $k$ based on the discrepancy between $\beta_{j,k}$ and all other coefficients, $(\beta_{j,l})_{l\neq k}$. That is, for each index $k$, we choose the words $j$ with the largest values of $\max\{\min\{\beta_{j,k}-\beta_{j,l}:l\neq k\},0\}$. We display the representative words for each latent variable in the bottom row of Figure 5. Here, each latent variable is named based on the representative words and the held-out newsgroup categories.
@@
Compared to the held-out tree structure of the 12 newsgroup labels (see Figure S.7 in the Supplement), one can see that the DDE has discovered a lower-dimensional structure in Figure 5. The latent structure in DDEs allow multiple parents for each topic and effectively model the complex label dependence. For example, ‘cars’, ‘software’, ‘location/names’ have both ‘recreation’ and ‘technology’ as parents, and many bottom-layer words are assigned to multiple topics. We also observe that similar true labels are combined into a single latent variable, for example ‘computer’ and ‘science’ are combined into ‘technology’ in the second latent layer, and ‘baseball’ and ‘hockey’ are combined as ‘sports’ in the bottom latent layer.
@@
We also compare our model fit to existing directed graphical models with matching latent dimensions: LDA (Blei et al.,, 2003) and DPFA-SBN (Gan et al.,, 2015). LDA has a single latent layer with mixed membership scores as continuous latent variables, and DPFA-SBN is a multilayer model with binary latent variables similar to DDEs. We consider the following three metrics widely used in topic modeling to measure different aspects of fit (Chen et al.,, 2023). The first measure is the perplexity, measuring the predictive likelihood of the words in the held-out set. While there are multiple notions of perplexities, we choose the definition that was adopted in similar Poisson-based models (Zhou et al.,, 2012; Gan et al.,, 2015; Ranganath et al.,, 2015, see Supplementary Material S.5.3 for more details). In addition to the train perplexity, we also display the test perplexity. The second measure is the average negative coherence, measuring the quality within each topic by computing $-\frac{1}{K^{(1)}}\sum_{k=1}^{K^{(1)}}\sum_{v_{1},v_{2}\in V_{k}}\log\left({(\text{freq}(v_{1},v_{2})+1)}/{\text{freq}(v_{2})}\right),$ where $V_{k}$ is the top $15$ representative words for the $k$th topic, and the function “freq” counts the number of documents containing the words specified in the input argument. The third measure is the similarity, computing the number of overlapping words within representative words across different topics: $\sum_{1\leq k_{1}<k_{2}\leq K^{(1)}}\sum_{v_{1}\in V_{k_{1}},v_{2}\in V_{k_{2}}}I(v_{1}=v_{2}).$ For all three measures, smaller values are better.
@@
Table 4 summarizes the results and shows the promising fit of DDE. Compared to other models with the same latent dimensions, DDEs have better test perplexity and similarity. The similarity measure shows that while the other model fits exhibit common representative words among different topics, the representative words learned from DDE are entirely disjoint and effectively represent different topics. In terms of coherence, the DDE fit is better than LDA but worse than DPFA. We have also fit models with larger dimensions for comparison, by considering LDA with $256$ latent variables, and DPFA with $K^{(1)}=128,K^{(2)}=64$ latent variables. The dimension for LDA is motivated by that the DDE has $2562^{8}=$ 256 mixture components; while for DPFA, this is the same latent dimension specified in the original paper (Gan et al.,, 2015). We can see that while considering a larger latent dimension may help in terms of perplexity, this leads to a loss of the within-topic coherence as well as dilutes the boundary of each topic, and hence gives less interpretable results.
@@
6.3 Continuous Data: Lognormal-DDE for Response Times
We apply the DDE to a response time dataset from the Trends in International Mathematics and Science Study (TIMSS) (Fishbein et al.,, 2021). We analyze the eighth-grade students’ response time for an internet-based mathematics assessment. After preprocessing, the dataset consists of $N=526$ students’ response times for $J=29$ questions. Here, we fit the two-latent-layer DDE where a lognormal distribution is used to model the continuous positive response times (Lognormal-DDE): $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{lognormal}\big{(}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)},\gamma_{j}\big{)}$. This dataset includes additional information regarding the latent structure $\mathbf{G}^{(1)}$, which is the so-called $Q$-matrix in the cognitive diagnostic modeling literature (von Davier,, 2008; Lee and Gu,, 2024). The provisional $\mathbf{G}^{(1)}$ matrix specifies $K^{(1)}=7$ latent cognitive skills (Number, Algebra, Geometry, Data and Probability, Knowing, Applying, and Reasoning) as well as the skills that are required to solve each item. That is, $g^{(1)}_{j,k}=1$ if item $j$ requires latent skill $k$ to solve it. Hence, following the confirmatory latent variable modeling convention in psychometrics, we estimate the DDE parameters by fixing $\mathbf{G}^{(1)}$ to this given structure.

### v2

@@ 5 Simulation Studies
We conduct extensive simulation studies in various settings to assess the performance of the proposed computation pipeline (Algorithms 1 and 2) and validate our identifiability conditions (in Section 3.1).
Two-latent-layer DDEs with general response types.
First, we generate data from two-latent-layer DDEs, exploring a total of 90 true settings by varying the following:
@@
    two parameter values: see Supplement S.4.1,
@@
Here, we consider two sets of parameter values that each satisfy the strict and generic identifiability conditions in Theorems 1 and 2. Regarding the parameter dimensions, given a value of $K^{(2)}$, we set $K^{(1)}=3K^{(2)}$ and $J=9K^{(2)}$. This allows a large latent dimension with as many as $K^{(1)}=30$ binary latent variables in the shallowest latent layer.
@@
Figure 4: Estimation error for $\mathcal{G}$ and $\boldsymbol{\Theta}$ under the two-latent-layer DDE with strictly identifiable true parameters and various observed-layer parametric families.
@@
For each scenario, we run $100$ independent simulations, and display the estimation results in Figure 4. The results under the generic identifiable parameters and computation times are included in Supplement S.4.3. We measure the estimation accuracy of the graphical matrices $\mathcal{G}$ by computing the average entrywise accuracy. For the continuous parameters $\boldsymbol{\Theta}$, we report the root mean squared error (RMSE). Under all response types and parameter values, the estimation errors for both $\mathcal{G}$ and $\boldsymbol{\Theta}$ decrease as the sample size $N$ increases. This empirically justifies the identifiability and consistency results. Additionally, by comparing the estimation accuracy across different parametric families, we observe that the Bernoulli is the most challenging to estimate, and the Normal is the easiest.
Simulation studies under deeper models.
We assess the scalability of our proposed method for deeper models (with $D=3,4,5$ latent layers) with potentially high-dimensional responses (with the observed data dimension set to $J=54,162,486$, respectively). For all settings, we consider Normal observations with $K^{(D)}=2$ latent variables for the top (deepest) layer and $K^{(d-1)}=3K^{(d)}$ variables for the $d$th layer for each $d=D,\ldots,2$.
We summarize the estimation accuracy of the graph structure in each layer and the average runtime for DDEs with $D=5$ latent layers in Table 1. The results for $D=3,4$ are displayed in Supplement S.4.6. The computation time illustrates that our proposed method is scalable to deeper models. More specifically, the $D=3,4$ case only took a few minutes, and the most challenging case with $D=5$ and $N=16,000$ also took less than an hour on a personal laptop. In terms of estimation accuracy of graph structures $\mathcal{G}$, the accuracy is higher for shallower layers than for deeper layers. This results from the accumulation of uncertainty for deeper layers, which is inevitable as each layer consists of stochastic latent variables. Note that even for such deeper models with $D=4,5$ latent layers, the shallower structures $\mathbf{G}^{(1)},\mathbf{G}^{(2)}$ are recovered with high accuracy. This indicates that one can include additional latent layers to increase the models’ representational power, without sacrificing the accuracy of learning graph structures closer to the data layer.
[…]
$\mathbf{G}^{(1)}$ 1.00 1.00 1.00 1.00 1.00 1.00
$\mathbf{G}^{(2)}$ 0.81 0.90 0.95 0.97 0.98 0.98
$\mathbf{G}^{(3)}$ 0.77 0.83 0.86 0.88 0.90 0.92
$\mathbf{G}^{(4)}$ 0.61 0.62 0.65 0.67 0.66 0.74
$\mathbf{G}^{(5)}$ 0.60 0.59 0.62 0.62 0.65 0.64
runtime (s) 220 259 326 563 893 2925
Table 1: Average entrywise-accuracy of estimating the graphical matrices and runtime in seconds for DDEs with $D=5$ latent layers.
Ablation studies.
As our model closely resemble DBNs, we mainly compare DDEs with DBNs. For fair comparison, we considered two settings: (a) data generated from a Berounlli-DDE, and (b) data generated from a DBN with identical coefficients. The results in Figure 5 demonstrate that our proposed algorithm has better estimation accuracy for both the graphical structure and the continuous parameters under both well-specified and mis-specified settings, even when data are generated from a DBN with undirected edges in the top layer. This is because the DBN algorithm does not learn sparse coefficients, and also fundamentally suffers from local optima due to a random initialization. Note that our identifiability results in Propositions S.1-2 immediately guarantee the identifiability of DBNs, and theoretically justify the above positive results of our algorithm.
The ablation studies also indicate that the current implementations of DDEs are slower than DBNs. We believe that our algorithm can be further scaled up by replacing the M-step in the SAEM algorithm to first-order optimization methods (e.g. gradient ascent or stochastic gradient ascent), which we leave for future work.
Figure 5: Comparison of Bernoulli-DDE vs DBN. True data are generated from a DDE in (a)–(b), and from a DBN in (c)–(d). Red and blue lines show results obtained by the DBN algorithm and our DDE algorithm, respectively. (a) and (c): estimation error for the graphs $\mathcal{G}$. (b) and (d): estimation error for continuous parameters $\Theta$. Smaller values are better.
Additional simulation results.
@@
In Supplement S.4.4 and S.4.6, we present additional simulation results where the latent dimension $\mathcal{K}$ is unknown. We illustrate that the spectral-gap estimator has near-perfect selection accuracy for a large $N$ (e.g., larger than $4000$), and is superior in terms of both accuracy and computation time compared to alternatives.
@@ 6.1 Binary Data: Bernoulli-DDE for MNIST Handwritten Digits
${\boldsymbol{\beta}}_{0,0}^{(1)}$ ${\boldsymbol{\beta}}_{0,1}^{(1)}$ ${\boldsymbol{\beta}}_{0,2}^{(1)}$ ${\boldsymbol{\beta}}_{0,3}^{(1)}$ ${\boldsymbol{\beta}}_{0,4}^{(1)}$ ${\boldsymbol{\beta}}_{0,5}^{(1)}$
Table 2: From left to right: Estimated basis images (reshaped from the estimated coefficients ${\boldsymbol{\beta}}_{0,k}^{(1)}$) from the MNIST data, which define sparse subregions in the $28\times 28$ image. The second and third row shows the most significant positive and negative parts of the basis images by thresholding at the value $\pm 1.5$.
We fit the two-latent-layer DDE with the latent dimensions set to $K^{(1)}=5$ and $K^{(2)}=2$ (see the Supplementary Material S.5.2 for the rationale for this choice). Table 2 displays the first-layer coefficients $\mathbf{B}^{(1)}$ by rearranging the each column into the original $28\times 28$ grid. Note that the value of $\mathbf{B}^{(1)}$ is in the logit-scale, so the negative coefficients make the corresponding pixel more likely to be zero. From the reshaped columns ${\boldsymbol{\beta}}_{0,0}^{(1)}$, ${\boldsymbol{\beta}}_{0,1}^{(1)}$, …, ${\boldsymbol{\beta}}_{0,5}^{(1)}$ of $\mathbf{B}^{(1)}$, we can interpret the meaning of each latent variable: $A_{1}^{(1)}=1$ indicates a zero-like shape, $A_{5}^{(1)}=1$ indicates symmetric curves on the upper-left and bottom-right corners, and the other latent variables represent different rotations. Additionally, using the deeper graphical matrix $\mathbf{G}^{(2)}$ displayed in the Supplementary Material S.5.4, we can also interpret the top layer latent variables as broader information about the images. For example, $A_{1}^{(2)}$ indicates large pixel density and $A_{2}^{(2)}$ indicates symmetry with respect to the x-axis.
@@
$(\widehat{\mathbf{A}}^{(1)}_{i},\widehat{\mathbf{A}}^{(2)}_{i})=\operatorname*{argmax}_{\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}},~\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K_{2}}}\varphi_{i,\boldsymbol{\alpha}^{(1)},\boldsymbol{\alpha}^{(2)}};\quad\boldsymbol{\alpha}^{(d)}\in\{0,1\}^{K^{(d)}},~d=1,2.$ (10)
Then, a decision tree that classifies the binary latent representations to the categorical label is built by using the misclassification error as the splitting criteria; see the left panel of Figure 6. The center and right panels display the estimated latent traits and estimated labels. Our classifier leads to a high train/test accuracy of $92.0\%/92.6\%$, even though our model is not fine-tuned for image classification. Although the state-of-art machine learning methods can achieve an accuracy as high as $97\%$ (Monnier et al., , 2020), we point out that the main goal here is not classification, but on interpretability and parsimony; indeed, the DDE uses more limited information and is a less complex but more interpretable model that provides the generative process for the image. In Table 4, we compare the classification accuracy and pixel-wise reconstruction accuracy of the two-latent-layer DDE with alternative interpretable models (latent class model (LCM) and the one-latent-layer DDE), as well as popular unsupervised machine learning models (two-latent-layer DBN and VAE). The results show that the two-latent-layer DDE performs well for both measures. We also display example images generated from DDE alongside their latent representations in Table 3, which illustrate various handwriting styles for each digit. We provide implementation details, detailed comparison with iVAEs, and additional visualization in Supplementary Material S.5.3.
$[1,0,1,0,1]$ $[0,0,0,0,0]$ $[0,1,0,1,1]$ $[0,0,1,1,1]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $[1,0,1,1,1]$ $[0,0,0,1,0]$ $[0,1,1,0,1]$ $[0,0,1,1,0]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Table 3: Example images generated from DDE and their latent representations. Accuracy LCM 1-DDE 2-DDE 2-DBN VAE Train classif. (%) 89.5 84.8 92.0 89.3 97.1 Test classif. (%) 90.9 86.2 92.6 90.6 92.0 Train recon. (%) 48.2 79.5 79.6 76.4 82.5 Test recon. (%) 48.5 79.1 79.9 77.0 82.7 Table 4: Classification accuracy and pixel-wise reconstruction accuracy for the MNIST dataset.
@@ 6.2 Count Data: Poisson-DDE Hierarchical Topic Modeling
Y1.0e0]ptrecreationcarssportsmotorcyclesoftwarelocation/nameshardwaregraphicsspacetechnologygeb bike banks riding pitt dod bmw gordonhockeynhlplayoffsteamsplayersleagueseasonbaseballenginecheappartsturnedmodelsdealermarketcarfilesdirectoryfilemscopydeletedsizecharlespeter ny chicagodevicesperiodyorkthomasusamotherboard dx CPU mhz controller board ram cardsmouse motif lines color display screen window imageorbit earth research science nasa development project institute
@@
We analyze the text corpus from the 20 newsgroups dataset (Lang, , 1995), which was previously analyzed by other topic models with binary latent variables (Srivastava et al., , 2013; Gan et al., , 2015). After preprocessing and focusing on 12 newsgroups, the dataset consists of $N=5,883$ documents and $J=653$ words. We fit the two-latent-layer DDE with a Poisson-distributed data layer (Poisson-DDE) $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{Poi}\big{(}\exp\big{[}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)}\big{]}\big{)}$ and latent dimensions $K^{(1)}=8,K^{(2)}=2$, and display the estimated latent structure in Figure 7. Additional details behind this choice are given in Supplementary Material S.5.2. To better interpret individual latent variables, we define representative words for each topic $k$ based on the discrepancy between $\beta_{j,k}$ and all other coefficients, $(\beta_{j,l})_{l\neq k}$. That is, for each index $k$, we choose the words $j$ with the largest values of $\max\{\min\{\beta_{j,k}-\beta_{j,l}:l\neq k\},0\}$. We display the representative words for each latent variable in the bottom row of Figure 7. Here, each latent variable is named based on the representative words and the held-out newsgroup categories.
@@
Compared to the held-out tree structure of the 12 newsgroup labels (see Figure S.7 in the Supplement), the DDE discovered a lower-dimensional structure in Figure 7. The latent structure in DDEs allow multiple parents for each topic and effectively model the complex label dependence. For example, ‘cars’, ‘software’, ‘location/names’ have both ‘recreation’ and ‘technology’ as parents, and many bottom-layer words are assigned to multiple topics. We also observe that similar true labels are combined into a single latent variable, for example ‘computer’ and ‘science’ are combined into ‘technology’ in the second latent layer, and ‘baseball’ and ‘hockey’ are combined as ‘sports’ in the bottom latent layer.
@@
We also compare our model fit to existing directed graphical models with matching latent dimensions: LDA (Blei et al., , 2003) and DPFA-SBN (Gan et al., , 2015). LDA has a single latent layer with mixed membership scores as continuous latent variables, and DPFA-SBN is a multilayer model with binary latent variables similar to DDEs. We consider the following three metrics widely used in topic modeling to measure different aspects of fit (Chen et al., , 2023). The first is the perplexity, measuring the predictive likelihood of the words in the held-out set. The second is the average negative coherence, measuring the quality within each topic by computing $\frac{1}{K^{(1)}}\sum_{k=1}^{K^{(1)}}\sum_{v_{1},v_{2}\in V_{k}}\log\left({(\text{freq}(v_{1},v_{2})+1)}/{\text{freq}(v_{2})}\right),$ where $V_{k}$ is the top $15$ representative words for the $k$th topic, and the function “freq” counts the number of documents containing the words specified in the input argument. The third is the similarity, computing the number of overlapping representative words across different topics: $\sum_{1\leq k_{1}<k_{2}\leq K^{(1)}}\sum_{v_{1}\in V_{k_{1}},v_{2}\in V_{k_{2}}}I(v_{1}=v_{2}).$ For all three measures, smaller values are better.
@@
Table 5 summarizes the results and shows the promising fit of DDE. Compared to other models with the same latent dimensions, DDEs have better test perplexity and similarity. The similarity measure shows that while the other model fits exhibit common representative words among different topics, the representative words learned from DDE are entirely disjoint and effectively represent different topics. In terms of coherence, the DDE fit is better than LDA but worse than DPFA. We have also fit models with larger dimensions for comparison, by considering LDA with $256$ latent variables, and DPFA with $K^{(1)}=128,K^{(2)}=64$ latent variables. The dimension for LDA is motivated by that the DDE has $2^{8}=256$ mixture components; while for DPFA, this is the same latent dimension specified in the original paper (Gan et al., , 2015). We can see that while considering a larger latent dimension may help in terms of perplexity, this leads to a loss of the within-topic coherence as well as dilutes the boundary of each topic, and hence gives less interpretable results.
@@
6.3 Multimodal Educational Data: Bernoulli-Lognormal-DDE
We apply the DDE to an educational assessment dataset from the Trends in International Mathematics and Science Study (TIMSS) (Fishbein et al., , 2021). We analyze the eighth-grade students’ responses for an internet-based mathematics assessment. As the assessment is electronically conducted, multiple modalities of information are recorded. Here, we focus on two important modalities: binary response accuracy and continuous response time. For each individual student, our data consists of response accuracy (whether the student gave a correct answer) and response time (how long the student took) for each of the $J/2=29$ items. We use the same latent variables to model both data modalities and model the binary response accuracy via Bernoulli distributions, and model the continuous positive response times via Lognormal distributions: $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{lognormal}\big{(}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)},\gamma_{j}\big{)}$.